Jack is 2 times older than his sister and his father is 25 years older than Jack. All of their ages put together are 53. How old is Jack and his father and sister
Answer by ikleyn(52800)

.
Jack is 2 times older than his sister and his father is 25 years older than Jack.
All of their ages put together are 53. How old is Jack and his father and sister
~~~~~~~~~~~~~~~~~~~~~~~

Sister + Jack + Father = 53,

S + 3S + (3S+25) = 53,      (2 times older = 3 times as old as . . . )

7S + 25 = 53,

7S = 53 - 25 = 28.


The sister is 4 years old, Jack is 12 years old, the father is 12 + 25 = 37 years old

Check.  4 + 12 + 37 = 53.   Correct !